# On the use of the Gram matrix for multivariate functional principal components analysis

### v1

@@ 5.1 Simulation experiments
We consider two simulation scenarios. One consists of multivariate functional data with univariate features defined on one-dimensional domains and the other consists of univariate functional data defined on a two-dimensional domain.
Scenario 1. The simulation setting is based on the simulation in Happ and Greven, (2018). The data-generating process is based on a truncated version of the Karhunen-Loève decomposition. First, we generate a large orthonormal basis $\{\psi_{k}\}_{1\leq k\leq K}$ of $\mathcal{L}^{2}\left(\mathcal{T}\right)$ on an interval $\mathcal{T}=[0,T]\subset\mathbb{R}$. We fix $T_{1}=0$ and $T_{P+1}=T$ and we generate $P-1$ cutting points $T_{2},\dots,T_{P}$ uniformly in $\mathcal{T}$ such that $0=T_{1}<\cdots<T_{P}<T_{P+1}=T$. Let $s_{1},\dots,s_{P}\in\{-1,1\}$ be coefficients that randomly flip the eigenfunctions with probability $0.5$. The univariate components of the eigenfunctions are then defined as
$\phi_{k}^{(p)}(t_{p})=s_{p}{\left.\kern-1.2pt\psi_{k}\vphantom{\big{|}}\right|_{[T_{p},T_{p+1}]}}\left(\frac{t_{p}-T_{p}}{T_{p+1}-T_{p}}\right),\quad k=1,\dots,K,\quad p=1,\dots,P.$ (42)
The notation ${\left.\kern-1.2pt\phi_{k}\vphantom{\big{|}}\right|_{[T_{p},T_{p+1}]}}$ is the restriction of the function $\phi_{k}$ to the set $[T_{p},T_{p+1}]$. The set of multivariate functions $\{\psi_{k}\}_{1\leq k\leq K}$ is an orthonormal system in $\mathcal{H}\coloneqq\mathcal{L}^{2}\left(\mathcal{T}_{1}\right)\times\dots\times\mathcal{L}^{2}\left(\mathcal{T}_{P}\right)$ with $\mathcal{T}_{p}=[0,1]$. Each curve is then simulated using the truncated multivariate Karhunen-Loève expansion (31):
$X(\mathbf{t})=\sum_{k=1}^{K}\mathfrak{c}_{k}\phi_{k}(\mathbf{t}),\quad\mathbf{t}\in\mathcal{T},$ (43)
where the scores $\mathfrak{c}_{k}$ are sampled as random normal variables with mean 00 and variance $\lambda_{k}$. The eigenvalues $\lambda_{k}$ are defined with an exponential decrease, $\lambda_{k}=\exp(-(k+1)/2)$. We simulate, for each replication of the simulation, $N=25,50,75$ and $100$ observations. Similarly, each component is sampled on a regular grid of $M=25,50,75$ and $100$ sampling points. We compare the methods for $P=2,10,20$ and $50$ features and we set $K=10$.
@@
Scenario 2. The data generating process is again based on a truncated version of the Karhunen-Loève decomposition. First, we generate an orthonormal basis $\{\phi_{k}\}_{1\leq k\leq K}$ of $\mathcal{L}^{2}\left(\mathcal{T}\right)$ on an interval $\mathcal{T}=[0,1]\times[0,1]$ as the tensor product of the first Fourier basis functions:
$\phi_{k}(s,t)=\psi_{l}(s)\otimes\psi_{m}(t),\quad s,t\in[0,1],\quad k=1,\dots,K,$ (44)
@@
where $\psi_{l}$ and $\psi_{m}$ are elements of the Fourier basis. Each curve is then simulated using the truncated multivariate Karhunen-Loève expansion (31):
$X(s,t)=\sum_{k=1}^{K}\mathfrak{c}_{k}\phi_{k}(s,t),\quad s,t\in[0,1],$ (45)
where the scores $\mathfrak{c}_{k}$ are defined as for the Scenario 1. We simulate, for each replication of the simulations, $N=25,50,75$ and $100$ observations. Similarly, each component is sampled on a regular grid of $M=25\times 25,50\times 50,75\times 75$ and $100\times 100$ sampling points. We set $K=10$.
@@ 5.2 Simulation results
We compared MFPCA using the diagonalization of the covariance operator and using the diagonalization of the Gram matrix in terms of their CT, estimation of eigenvalues, estimation of eigenfunctions, and reconstruction of curves. We fix the number of retained components to be $5$ for each simulation of both scenarios. Each experiment is repeated $500$ times. The results are presented below.
@@
Computational time. To compare the computational time of the diagonalization of the covariance operator and the diagonalization of the Gram matrix, we measured the time it took for each method to complete the MFPCA for each simulated dataset. Figure 5 shows the kernel density estimates of the ratio of CT for each method across all sample sizes, number of sampling points and number of features. For Scenario 1, we found that the diagonalization of the covariance has a shorter CT compared to the diagonalization of the Gram matrix for most combinations of sample sizes, number of functions and number of sampling points. It is faster to use the Gram matrix if the number of observations is low compared to the number of sampling points.
Figure 5: Ratio of computation time for Scenario 1 between the Gram matrix method and the covariance operator method. Each univariate component is defined on a one-dimensional domain. $N$ is the number of observations, $M$ is the number of sampling points per curve and $P$ is the number of features.
Figure 6 shows the kernel density estimates of the ratio of CT for each method across all sample sizes and number of sampling points. For Scenario 2, we found that the diagonalization of the Gram matrix has a shorter CT compared to the diagonalization of the covariance operator across all sample sizes and number of sampling points.
@@
Figure 6: Ratio of computation time for Scenario 2 between the Gram matrix method and the covariance operator method. $N$ is the number of observations and $M\times M$ is the number of sampling points per images.
@@
The shorter computational time of the diagonalization of the Gram matrix for Scenario 2 makes it a more efficient option for analyzing two and higher-dimensional functional datasets. It is however worth noting that the computational time can still vary depending on the specific implementation of each method, the computational resources available, and the complexity of the dataset (number of observations, number of sampling points, etc.).
@@
Eigenvalues estimation. To compare the estimation of the eigenvalues between the diagonalization of the covariance operator and the diagonalization of the Gram matrix, we calculated the ratio of the $\log-\text{AE}$ (39) between the estimated eigenvalues and the true eigenvalues for each simulated dataset and for the first five eigenvalues. Figure 7 shows the boxplots of the $\log-\text{AE}$ for each method across all sample sizes, number of sampling points and number of features for Scenario 1. We found that the two methods behave similarly for all considered settings.
Figure 7: Ratio of $\log-$AE for Scenario 1 between the Gram matrix method and the covariance operator method. Each univariate component is defined on a one-dimensional domain. $N$ is the number of observations, $M$ is the number of sampling points per curve and $P$ is the number of features.
Figure 8 shows the boxplots of the $\log-\text{AE}$ for each method across all sample sizes and number of sampling points for Scenario 2. We found that the FCP-TPA gives slighty better estimation of the eigenvalues.
Figure 8: Ratio of $\log-$AE for Scenario 2 between the Gram matrix method and the covariance operator method. $N$ is the number of observations and $M\times M$ is the number of sampling points per images.
Eigenfunctions estimation. To compare the estimation of the eigenfunctions between the diagonalization of the covariance operator and the diagonalization of the Gram matrix, we calculated the ratio of the ISE (38) between the estimated eigenfunctions and the true eigenfunctions for each simulated dataset and for the first five eigenfunctions. Figure 9 shows the boxplots of the ISE for each method across all sample sizes, number of sampling points and number of features for Scenario 1. We found that the two methods behave similarly for all considered settings. For $P=10,20$ and $50$, the results are identical.
Figure 9: Ratio of ISE for Scenario 1 between the Gram matrix method and the covariance operator method. Each univariate component is defined on a one-dimensional domain. $N$ is the number of observations, $M$ is the number of sampling points per curve and $P$ is the number of features.
Figure 10 shows the boxplots of the ISE for each method across all sample sizes and number of sampling points for Scenario 2. We found that the decomposition of the Gram matrix gives better estimation of the eigenfunctions compared to the FCP-TPA, especially when the number of observations increases.
Figure 10: Ratio of ISE for Scenario 2 between the Gram matrix method and the covariance operator method. $N$ is the number of observations and $M\times M$ is the number of sampling points per images.
Curves reconstruction. To compare the quality of the reconstruction of the curves between the diagonalization of the covariance operator and the diagaonlisation of the Gram matrix, we calculated the ratio of the MISE (40) between the reconstruction of the curves and the true curves for each simulated dataset. Figure 11 shows the boxplots of the ISE for each method across all sample sizes, number of sampling points and number of features for Scenario 1. We found that the diagonalization of the Gram matrix gives slightly better results for all considered settings.
Figure 11: Ratio of MISE for Scenario 1 between the Gram matrix method and the covariance operator method. Each univariate component is defined on a one-dimensional domain. $N$ is the number of observations, $M$ is the number of sampling points per curve and $P$ is the number of features.
Figure 12 shows the boxplots of the ISE for each method across all sample sizes and number of sampling points for Scenario 2. Similarly to Scenario 1, we found that the decomposition of the Gram matrix gives better estimation of the true curves compared to the FCP-TPA, especially when the number of observations increases.
Figure 12: Ratio of MISE for Scenario 2 between the Gram matrix method and the covariance operator method. $N$ is the number of observations and $M\times M$ is the number of sampling points per images.
@@ 6 Discussion and conclusion
MFPCA is a fundamental statistical tool for the analysis of multivariate functional data, which enables us to capture the variability in observations defined by multiple curves. In this paper, we have described the duality between rows and columns of a data matrix within the context of multivariate functional data. We have proposed to use this duality to estimate the eigencomponents of the covariance operator in multivariate functional datasets. By comparing the results of the two methods, we provide the researcher with guidelines for determining the most appropriate method within a range of functional data frameworks. Overall, our simulations showed that the diagonalization of the covariance operator or of the Gram matrix gave similar results in terms of the estimation of the eigenvalues, the eigenfunctions and reconstruction of the curves for one-dimensional multivariate functional data, while the performance of the diagonalization of the Gram matrix outperformed the FCP-TPA for higher-dimensional functional datasets. Regarding the computation time, the use of the covariance operator is faster in most cases for multivariate functional data defined on unidimensional domains. The only situation where the use of the Gram matrix is quicker is when $M\gg N$. For data defined on higher-dimensional domains, diagonalization of the Gram matrix is faster. In conclusion, we recommend to use the covariance operator for multivariate functional data with all features defined on a one-dimensional domeins (curves) and for a number of observations larger or comparable to the number of sampling points regardless of the number of features. If the data are defined on multi-dimensional domains (images) or the number of sampling points is much higher than the number of observations, we advise using the Gram matrix.
@@
Utilizing the Gram matrix enables the estimation of the number of components retained via the percentage of variance explained by the multivariate functional data, whereas the decomposition of the covariance operator necessitates the specification of the percentage of variance accounted for by each individual univariate feature. Specifying the percentage of variance explained for each feature does not guarantee that we recover the nominal percentage of variance explained for the multivariate data. Although we have not investigated the extent to which this might be important, the duality relation derived in this work provides a direct solution to the problem. Future work could investigate the settings in which univariate variance-explained cutoffs fail to retain the correct percentage of variance explained in multivariate functional data, and hence where the Gram matrix approach may be preferred.
In practice, observations of (multivariate) functional data are often subject to noise. As we recommend the use of the Gram matrix solely for densely sampled functional datasets, individual curve smoothing should suffice to approximate the Gram matrix in such cases. The estimation of the Gram matrix in the context of sparsely sampled functional data is however deemed irrelevant, given our findings that the utilization of the covariance operator for the estimation of the eigencomponents yields comparable results, while typically requiring less computational time.
@@
The open-source implementation can be accessed at https://github.com/StevenGolovkine/FDApy, while scripts to reproduce the simulations are at https://github.com/FAST-ULxNUIG/geom_mfpca.

### v2

@@ 5.1 Simulation experiments
The simulation setting is based on the Setting 3 of the simulation in Happ and Greven, (2018). The data generating process is based on a truncated version of the Karhunen-Loève decomposition. We simulate multivariate functional data with $P=2$ components. For the first component, we generate an orthonormal basis $\{\phi^{(1)}_{k}\}_{1\leq k\leq K}$ of $\mathcal{L}^{2}\left(\mathcal{T}_{1}\right)$ on an interval $\mathcal{T}_{1}=[0,1]\times[0,0.5]$ as the tensor product of the first Fourier basis functions:
$\phi^{(1)}_{k}(s,t)=\psi_{l}(s)\otimes\psi_{m}(t),\quad s\in[0,1]\text{ and }t\in[0,0.5],\quad k=1,\dots,K,$ (44)
@@
where $\psi_{l}$ and $\psi_{m}$ are elements of the Fourier basis. For the second component, we generate an orthonormal Legendre basis $\{\phi^{(2)}_{k}\}_{1\leq k\leq K}$ of $\mathcal{L}^{2}\left(\mathcal{T}_{2}\right)$ on an interval $\mathcal{T}_{2}=[-1,1]$. To ensure orthonormality, the basis $\phi^{(1)}_{k}$ and $\phi^{(2)}_{k}$ are weighted by random factors $\alpha^{1/2}$ and $(1-\alpha)^{1/2}$, respectively, where $\alpha\sim\mathcal{U}(0.2,0.8)$. Each curve is then simulated using the truncated multivariate Karhunen-Loève expansion (34):
$X(\mathbf{t})=\sum_{k=1}^{K}\mathfrak{c}_{k}\phi_{k}(\mathbf{t}),\quad\mathbf{t}\in\mathcal{T}\coloneqq\mathcal{T}_{1}\times\mathcal{T}_{2},$ (45)
where $\phi_{k}=(\phi^{(1)}_{k},\phi^{(2)}_{k})^{\top}$ and the scores $\mathfrak{c}_{k}$ are sampled as random normal variables with mean 00 and variance $\lambda_{k}$. The eigenvalues $\lambda_{k}$ are defined with an exponential decrease, $\lambda_{k}=\exp(-(k+1)/2)$ We simulate, for each replication of the simulation, $N=50,100$ and $250$ observations. The first component is sampled on a regular grid of $M^{(1)}=11\times 11,26\times 26$ and $101\times 51$ sampling points. The second component is sampled on a regular grid of $M^{(2)}=21,51$ and $201$ sampling points. We set $K=25$. We also consider data with measurement errors. In that case, we observe data from (11), where $\sigma^{2}=(0.25,0.25)$. Finally, we study sparse data as in Happ and Greven, (2018), with medium ($50\%-70\%$ missing) and high ($90\%-95\%$ missing) sparsity in both components. For this setting, no noise is added to the data.
The (Tensor) PCA method is based on the univariate estimation of $K^{(1)}=20$ eigenimages and $K^{(2)}=15$ eigenfunctions. The estimation of the eigenimages is performed with the FCP-TPA where the smoothing parameters are chosen via cross-validation in $[10^{-5},10^{5}]$ for dense, sparse and noisy data. In the case of sparse data, we also linearly interpolate the two-dimensional observations to get data that are regularly sampled to run the FCP-TPA. The estimation of the eigenfunctions is done using the PACE algorithm (Yao et al.,, 2005) with P-splines to smooth the mean and covariance functions. Penalties are chosen using cross-validation (Eilers and Marx,, 1996).
For the 2D/1D B-splines method, $X^{(1)}$ is expanded in tensor products of $K^{(1)}=13\times 13$ B-splines and $X^{(2)}$ is expanded in $K^{(2)}=13$ B-splines. For dense data, no penalty is involved, while for sparse and noisy data, penalties are chosen using cross-validation.
Finally, the Gram method is based on the observed data points in the dense case, on the linear interpolation of the data in the sparse case (Benko et al.,, 2009) and on the P-splines smoothing of the data in the noisy case.
@@ 5.2 Simulation results
We compared MFPCA results of the different methods in terms of their CT, estimation of eigenvalues (RSE), estimation of eigenfunctions (ISE), and reconstruction of curves (MRSE). We fix the number of retained components to be $12$ for each simulation scenario. Each experiment is repeated $200$ times. The results are presented below.
@@
Computational time. To compare the computational time of the different methods, we measured the time taken for each method to complete the MFPCA for each simulated dataset in the dense and noiseless case. Figure 5 shows the kernel density estimates of the ratio of CT for each method across all sample sizes and number of sampling points. Except when we have many observations with few sampling points, the Gram method is faster than the other two methods. This result is coherent with the computational complexity derived in Section 4.3.
@@
Figure 5: Ratio of computation time in the dense case between the (Tensor) PCA and 2D/1D B-Splines methods and the Gram method. $N$ is the number of observations, $M$ is the number of sampling points per curve (the first two numbers are for the images and the last one is for the curves).
@@
The shorter computational time of the diagonalization of the Gram matrix makes it a more efficient option for analyzing two and higher-dimensional functional datasets as the number of sampling points tend to grow faster in these situations. However, the computational time can still vary depending on the specific implementation of each method, the computational resources available, the complexity of the dataset (number of observations, number of sampling points, etc.) and whether a smoothing step needs to be included.
@@
Eigenvalue estimation. To compare the estimation of the eigenvalues between the different methods, we calculated the RSE (42) between the estimated eigenvalues and the true eigenvalues for each simulated dataset and for the twelve estimated eigenvalues. Figure 6 shows the boxplots of the RSE for each method across all sample sizes and number of sampling points. We found that the three methods behave similarly for the settings with a moderately large to a large number of sampling points. When we observe only a few sampling points, $M^{(1)}=11\times 11$ and $M^{(2)}=21$, the quality of the estimation is approximately the same for the first three eigenvalues for all methods, but from the fourth eigenvalues, the 2D/1D B-splines and the Gram methods give slightly better results than the (Tensor) PCA method. The results for the sparse and noisy cases are similar to the dense and noiseless case and are provided in the Appendix C.1.
Figure 6: RSE for the estimated eigenvalues for each method in the dense case. $N$ is the number of observations, $M$ is the number of sampling points per curve (the first two numbers are for the images and the last one is for the curves).
Eigenfunction estimation. To compare the estimation of the eigenfunctions between the different methods, we calculated the ISE (41) between the estimated eigenfunctions and the true eigenfunctions for each simulated dataset and for the twelve estimated eigenfunctions. Figure 7 shows the boxplots of the ISE for each method across all sample sizes and number of sampling points. The results are similar to those found when estimating the eigenvalues. When we observe only a few sampling points, the quality of the estimation of the eigenfunctions starts to decrease after the third component for the (Tensor) PCA method, while this is not the case for the other methods. The results for the sparse and noisy cases are similar to the dense and noiseless case and are provided in the Appendix C.1.
Figure 7: ISE for the estimated eigenfunctions for each method in the dense case. $N$ is the number of observations, $M$ is the number of sampling points per curve (the first two numbers are for the images and the last one is for the curves).
Curve reconstruction. To compare the quality of the reconstruction of the curves between the different methods, we calculated the MRSE (43) between the reconstruction of the curves and the true curves for each simulated dataset. Figure 8 shows the boxplots of the MRSE for each method across all sample sizes and number of sampling points. Except in the case of few sampling points where the (Tensor) PCA method does not perform well (because of the poor estimation of the eigenvalues and eigenfunctions), all methods give nearly the same results for each setting. The results for the sparse and noisy cases are similar and are provided in the Appendix C.1.
Figure 8: MRSE for the reconstructed curves for each method in the dense case. $N$ is the number of observations, $M$ is the number of sampling points per curve (the first two numbers are for the images and the last one is for the curves).
6 Application
We apply our methodology utilizing data to the National Basketball Association (NBA). To access comprehensive game data, we utilize the Python package nba_api [^1] https://github.com/swar/nba_api/, which provides access to the APIs of nba.com. The dataset encompasses shot location data from all NBA games spanning the seasons between $2018-2019$ and $2022-2023$. Filtering the dataset, we focus solely on players who have made more than $1000$ shots during this five-season period, resulting in a cohort of $131$ players. These players accounted for a total of $493723$ shots attempted, of which $234941$ ($47\%$) were successful. Subsequently, we exclude shots deemed impossible (e.g., out-of-bounds), leaving us with a dataset comprising $492621$ shots (see Figure 9 for the shots chart of Stephen Curry). We remove all the shots that are close to the hoop (a square of $2.7\times 2.8$ meters around the hoop), as these shots will be the most common and their pattern is not interesting from a shooting behavior perspective, and players that have made fewer than $100$ shots during the five-season period. This results in a cohort of $119$ players, with $186621$ attempted shots and $71893$ ($38.5\%$) made shots.
To analyze shooting behavior, we employ a 2-dimensional kernel density estimation for both the attempted and made shots, utilizing Silverman’s rule (Silverman,, 1986) for bandwidth estimation. The density estimation is conducted on a regularly spaced grid consisting of $201\times 201$ points (see Figure 9 for the estimated densities for Stephen Curry). We obtain observations of bivariate functional data (i.e., $P=2$ functional features), where both features are defined on a two-dimensional rectangular domain, and observed on an equidistant grid of size $201\times 201$.
Figure 9: Made/Missed shots chart (left) and the estimated densities (right) for Stephen Curry.
We estimate the functional principal components using the Gram, (Tensor) PCA and 2D/1D B-Splines methods. For the Gram method, we directly use the estimated densities to estimate the eigencomponents. For the (Tensor) PCA method, the estimation of the multivariate eigenimages is based on the univariate estimation of $30$ univariate eigenimages. The estimation of the (univariate) eigenimages is performed with the FCP-TPA where the smoothing parameters are chosen via cross-validation in $[10^{-5},10^{5}]$. For the 2D/1D B-splines method, the two components are expanded in tensor products of $13\times 13$ B-splines. We do not add a smoothing penalty in that case, as the smoothing step is performed in the kernel density estimation.
Figure 10 shows the estimated mean surfaces and eigenfunctions for the Gram method. The functional principal components, as a representation of the deviation from the mean surface function, may take negative values, while densities can only take positive values. For presentation purposes, the functional principal components have been normalized between $-1$ and $1$. Blue areas correspond to negative values of the functional components, and thus contribute negatively to the density, while red area corresponds to positive values of the functional components, and thus contribute positively to the density. The decomposition of the density of made shots is similar that of attempted shots. The obtained functional components can be explained as different shooting behavior. The first component, accounting for $56.9\%$ of the variance explained, contrasts two-point and three-point shooting. A player with a positive score for $\phi_{1}$ will tend to take/attempt shots more behind the three-point line, while if he has a negative score, he will prefer shooting within the two-points zone. For the second component ($11.5\%$ of the variance explained), the contrast is between the left and right sides of the court. Similarly, if a player has a positive (resp. negative) score for $\phi_{2}$, he will prefer to shoot on the right (resp. left) side of the court while looking at the hoop. The third component ($9.4\%$ of the variance explained) contrasts shooting from the wing with shooting in the axis of the hoop. Positive (resp. negative) score players will shot in the axis of the hoop (resp. from the wing). The fourth component ($4.7\%$ of the variance explained) is more difficult to explain from a shooting behavior perspective. It may be linked to players having a preferred shooting location so that most of their shots are taken from the same position. The estimation of the eigencomponents with the (Tensor) PCA and 2D/1D B-Splines methods are similar to the results with the Gram method and are thus provided in the Appendix C.
Figure 10: The estimated mean surfaces (first column) and the estimated eigenfunctions (second to fifth columns) for the shots dataset using the Gram method.
@@ 7 Discussion and conclusion
MFPCA is a fundamental statistical tool for the analysis of multivariate functional data, which enables us to capture the variability in observations defined by multiple curves. In this paper, we have described the duality between rows and columns of a data matrix within the context of multivariate functional data. We have proposed to use this duality to estimate the eigencomponents of the covariance operator in multivariate functional datasets. By comparing the results of the three methods, we provide the researcher with guidelines for determining the most appropriate method within a range of functional data frameworks. Overall, our simulations showed that the (Tensor) PCA, 2D/1D B-Splines and Gram methods give similar results in terms of the estimation of the eigenvalues, the eigenfunctions and reconstruction of the curves. Regarding the computation time, the use of the Gram matrix is faster in most cases. The only situation where the diagonalization of the covariance operator is quicker is when the number of observations is larger than the number of sampling points. In conclusion, regarding the reconstruction error and computational complexity, if the data are defined on multi-dimensional domains (images) or the number of sampling points is much higher than the number of observations, we advise using the Gram matrix. Another advantage of the Gram method is that it does not require basis expansion, such as the 2D/1D B-splines method, or estimate smoothing parameters, such as the (Tensor) PCA method when the data are observed without noise.
In practice, observations of (multivariate) functional data are often subject to noise. In this case, the diagonalization of the covariance operator seems preferable as the curves as to be smoothed beforehand to estimate the Gram matrix and yield similar results, except when the number of sampling is small where the Gram method is preferable. In the case of sparsely sampled functional data, we would advise to use the Gram method, using linear interpolation to estimate the inner-product matrix, as it does not require hyper-parameters estimation and gives slightly better results than the other methods.
@@
Utilizing the Gram matrix enables the estimation of the number of components retained via the percentage of variance explained by the multivariate functional data, whereas the decomposition of the covariance operator necessitates the specification of the percentage of variance accounted for by each individual univariate feature. Specifying the percentage of variance explained for each feature does not guarantee recovery of the nominal percentage of variance explained for the multivariate functional data (Golovkine et al.,, 2023). Although we have not investigated the extent to which this might be important, the duality relation derived in this work provides a direct solution to the problem. In settings where the univariate variance-explained cutoffs fail to retain the correct percentage of variance explained in multivariate functional data, the Gram matrix approach may be preferred.
@@
The open-source implementation can be accessed at https://github.com/StevenGolovkine/FDApy, while scripts to reproduce the simulations and data analysis are at https://github.com/FAST-ULxNUIG/geom_mfpca.